OpenStudy (anonymous):

(step by step) please explain to me how to complete square for example: x^2+3x+9=0 and please give more examples and explain. thanks :)

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=xGOQYTo9AKY

OpenStudy (mimi_x3):

x^2+ bx = (x+b/2)^2 - (b/2)^2 Substitute it in xD

OpenStudy (anonymous):

Possible intermediate steps: x^2+3 x+9 = 0 Subtract 9 from both sides: x^2+3 x = -9 Add 9/4 to both sides: x^2+3 x+9/4 = -27/4 Factor the left hand side: (x+3/2)^2 = -27/4 Take the square root of both sides: abs(x+3/2) = (3 i sqrt(3))/2 Eliminate the absolute value: x+3/2 = -(3 i sqrt(3))/2 or x+3/2 = (3 i sqrt(3))/2 Subtract 3/2 from both sides: x = -3/2 i (-i+sqrt(3)) or x+3/2 = (3 i sqrt(3))/2 Subtract 3/2 from both sides: x = -3/2 i (-i+sqrt(3)) or x = 3/2 i (i+sqrt(3))